(d^2y/dx2)-2(dy/dx)+2y=sin2x
1
Expert's answer
2020-10-01T12:05:13-0400

"\frac{d^2y}{dx^2}-2\frac{dy}{dx}+2y=\sin 2x"


Let us solve a characteristic equation:


"k^2-2k+2=0"

"(k-1)^2+1=0"

"(k-1)^2=-1"

"k_1=1+i, k_2=1-i"


The general solution is of the following form:

"y(x)=e^x(C_1\cos x+C_2\sin x)+y_p(x)," where "y_p(x)" is a particular solution, "C_1, C_2" are arbitrary constants.


The particular solution has the following form:

"y_p(x)=A\sin 2x+B\cos 2x", whrere "A" and "B" are constants.


"\frac{dy_p}{dx}=2A\cos 2x-2B\sin 2x"


"\frac{d^2y_p}{dx^2}=-4A\sin 2x-4B\cos 2x"


Put founded derivatives in the differential equation:


"-4A\sin 2x-4B\cos 2x-2(2A\cos 2x-2B\sin 2x)+2(A\sin 2x+B\cos 2x)=\sin 2x"


"(-4A+4B+2A)\sin 2x+(-4B-4A+2B)\cos 2x=\sin 2x"


"(-2A+4B)\sin 2x+(-4A-2B)\cos 2x=\sin 2x"


Equating the corresponding coefficients, we obtain a system of equations:


"\begin{cases} -2A+4B=1\\ -4A-2B=0 \end{cases}"


"\begin{cases} -2A+4B=1\\ B=-2A \end{cases}"


"\begin{cases} -10A=1\\ B=-2A \end{cases}"


"\begin{cases} A=-\frac{1}{10}\\ B=\frac{1}{5} \end{cases}"


Finally, the solution of the differential equation is the following:


"y(x)=e^x(C_1\cos x+C_2\sin x)-\frac{1}{10}\sin 2x+\frac{1}{5}\cos 2x."